OpenStudy (anonymous):

find the geometric mean of 2 and 8

OpenStudy (anonymous):

\[\sqrt{(2)(8)} = 4\]

OpenStudy (anonymous):

\sqrt(2*8) = 4

OpenStudy (anonymous):

thanks that made more sence